# Towards Principled Multi-Agent Task Agnostic Exploration

### v1

@@ Abstract
In reinforcement learning, we typically refer to task-agnostic exploration when we aim to explore the environment without access to the task specification a priori. In a single-agent setting the problem has been extensively studied and mostly understood. A popular approach cast the task-agnostic objective as maximizing the entropy of the state distribution induced by the agent’s policy, from which principles and methods follows. In contrast, little is known about task-agnostic exploration in multi-agent settings, which are ubiquitous in the real world. How should different agents explore in the presence of others? In this paper, we address this question through a generalization to multiple agents of the problem of maximizing the state distribution entropy. First, we investigate alternative formulations, highlighting respective positives and negatives. Then, we present a scalable, decentralized, trust-region policy search algorithm to address the problem in practical settings. Finally, we provide proof of concept experiments to both corroborate the theoretical findings and pave the way for task-agnostic exploration in challenging multi-agent settings.
Reinforcement Learning, Multi-Agent Reinforcement Learning, Task Agnostic Exploration, Unsupervised Reinforcement Learning
@@ 1 Introduction
Multi-Agent Reinforcement Learning (MARL, Albrecht et al., 2024) recently showed promising results in learning complex behaviors, such as coordination and teamwork (Samvelyan et al., 2019), strategic planning in the presence of imperfect knowledge (Perolat et al., 2022), and trading (Johanson et al., 2022). Just like in single agent RL, however, most of the efforts are focused on tabula rasa learning, that is, without exploiting any prior knowledge gathered from offline data and/or policy pre-training. Despite its generality, learning tabula rasa hinders MARL from addressing real-world situations, where training from scratch every time is slow, often expensive, and most of all unnecessary (Agarwal et al., 2022). In this regard, some progress has been made on techniques specific to the multi-agent setting, ranging from ad hoc teamwork (Mirsky et al., 2022) to zero-shot coordination (Hu et al., 2020).
@@
In single-agent RL, task-agnostic exploration, such as maximizing an entropy measure over the state space, was shown to be a useful tool for policy pre-training (Hazan et al., 2019; Mutti et al., 2021) and data collection for offline learning (Yarats et al., 2022). Recently, the potential of entropy objectives in MARL was empirically corroborated by a plethora of works (Liu et al., 2021a; Zhang et al., 2021; Xu et al., 2024) investigating entropic reward-shaping techniques to boost exploration. Yet, to the best of our knowledge, task-agnostic exploration has never been investigated for multi-agent scenarios explicitly, and, thus, the problem is far from being solved. Let us think of an illustrative example that highlights the central question of this work: multiple autonomous robots deployed in a collapsed building for a rescue operation mission. The robots’ main goal is to explore a large area to find and rescue injured humans, where exploration may involve coordinating with others to access otherwise inaccessible areas. Arguably, trying to enforce all robots to explore the entire area is inefficient and unnecessary. On the other hand, if everyone is focused on their own exploration, any incentive to collaborate with each other may disappear, especially when coordinating comes at a cost. Clearly, a third option is needed. Thus, many questions naturally arise: (i) How can task-agnostic exploration be defined in MARL? (ii) Are different formulations related in some way and when crucial differences emerge? (iii) How can we explicitly address multi-agent task-agnostic exploration in practical scenarios?
In this paper, we first provide a principled characterization of task-agnostic exploration in multi-agent settings by showing that the problem can take different, yet related, formulations. Specifically, the demarcation line is drawn by whether the agents are trying to jointly explore the space, or they neglect the presence of others and explore in a disjoint fashion, or, finally, whether they care of being able to explore the space together but as independent components of a mixture. We link these cases to three distinct objectives, each of them with specific pros and cons, and possibly leading to different behaviors. First, we formally show that these objectives are strictly related and enjoy similar behaviors in the ideal case of evaluating the agent’s performance over infinite realizations (or trials). Then, we shift the attention to the more practical scenario of reaching good performance over a handful, or even just one, trial. This is motivated by the fact that, even if we have access to a simulator to train our policies on many realizations, we often get to deploy them in a fair but limited amount of realizations. Interestingly, we show that different objectives enjoy rather different theoretical behaviors when optimized in such settings. Furthermore, we address the problem of how to optimize them, by introducing a decentralized multi-agent policy optimization algorithm, called Trust Region Pure Exploration (TRPE), explicitly addressing task-agnostic exploration over finite trials. Additionally, we test the algorithm on an illustrative proof of concept, showing its ability to optimize the objectives of interest, and, more importantly, showing how the diverse landscape of objectives turns out to be crucial for allowing effective coordinated exploration over short time-horizons. In particular, we show how optimizing for diverse exploration is not only crucial in practical scenarios, but is perhaps the only way to get relevant outcomes. We strengthen this claim by showing that this superiority is reflected in a higher effectiveness of policy pre-training for sparse-reward multi-agent tasks as well.
Contributions.
Throughout the paper, we make the following contributions:
  • We introduce a novel class of decision making problems, called Convex Markov Games, and use them to extend the task-agnostic exploration problem to multi-agent settings (Section 3);
  • We provide a theoretical characterization of the task-agnostic exploration problem, showing how the possible objectives are linked, and in what they are different (Section 4);
  • We design a decentralized trust-region policy search algorithm able to address the exploration task in its most practical formulation (Section 5);
  • We report empirical results that confirm the effectiveness of the algorithm in optimizing for their own objectives while highlighting the crucial differences of the different objectives. We showcase the limitations of more common objectives and the potentials of newly established mixture ones (Section 6).
@@ 2 Preliminaries
Notation.  In the following, we denote $[N]:=\{1,2,\ldots,N\}$ for a constant $N<\infty$. We denote a set with a calligraphic letter $\mathcal{A}$ and its size as $|\mathcal{A}|$. For a (finite) set $\mathcal{A}=\{1,2,\dots,i,\dots\}$, we denote with $-i=\mathcal{A}/\{i\}$ the set of all its elements out of the $i$-th one. We denote $\mathcal{A}^{T}:=\times_{t=1}^{T}\mathcal{A}$ the $T$-fold Cartesian product of $\mathcal{A}$. The simplex on $\mathcal{A}$ is denoted as $\Delta_{\mathcal{A}}:=\{p\in[0,1]^{|\mathcal{A}|}|\sum_{a\in\mathcal{A}}p(a)=1\}$ and $\Delta_{\mathcal{A}}^{\mathcal{B}}$ denotes the set of conditional distributions $p:\mathcal{A}\to\Delta_{\mathcal{B}}$. Let $X$ a random variable on the set of outcomes $\mathcal{X}$ and corresponding probability measure $p_{X}$, we denote the Shannon entropy of $X$ as $H(X)=-\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x))$. We denote $\mathbf{x}=(X_{1},\ldots,X_{T})$ a random vector of size $T$ and $\mathbf{x}[t]$ its entry at position $t\in[T]$.
Interaction Protocol.  As a base model for interaction, we consider finite-horizon Markov Games (MGs, Littman, 1994) without rewards. A MG $\mathcal{M}:=(\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},\mu,T)$ is composed of a set of agents $\mathcal{N}$, a set $\mathcal{S}=\times_{i\in[\mathcal{N}]}\mathcal{S}_{i}$ of states, and a set of (joint) actions $\mathcal{A}=\times_{i\in[\mathcal{N}]}\mathcal{A}_{i}$, which we let discrete and finite with size $|\mathcal{S}|,|\mathcal{A}|$ respectively. At the start of an episode, the initial state $s_{1}$ of $\mathcal{M}$ is drawn from an initial state distribution $\mu\in\Delta_{\mathcal{S}}$. Upon observing $s_{1}$, each agent takes action $a_{1}^{i}\in\mathcal{A}_{i}$, the system transitions to $s_{2}\sim\mathbb{P}(\cdot|s_{1},a_{1})$ according to the transition model $\mathbb{P}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$. The process is repeated until $s_{T}$ is reached and $s_{T}$ is generated, being $T<\infty$ the horizon of an episode. Each agent acts according to a policy, that can be either Markovian, i.e. $\pi^{i}\in\Delta_{\mathcal{S}}^{\mathcal{A}^{i}}$, or Non-Markovian over, i.e. $\pi^{i}\in\Delta_{\mathcal{S}^{t}\times\mathcal{A}^{t}}^{\mathcal{A}^{i}}$. [^1] In general, we will denote the set of valid per-agent policies with $\Pi^{i}$ and the set of joint policies with $\Pi$. Also, we will denote as decentralized policies the ones conditioned on either $\mathcal{S}_{i}$ or $\mathcal{S}^{t}_{i}\times\mathcal{A}^{t}_{i}$ for agent $i$, and centralized ones the one conditioned over the full state or state-actions sequences. It follows that the joint action is taken according to the joint policy $\Delta_{\mathcal{S}}^{\mathcal{A}}\ni\pi=(\pi^{i})_{i\in[\mathcal{N}]}$.
@@
Convex MDPs and Task-Agnostic Exploration.  Now, in the MDP setting ($N={1}$), the problem of task-agnostic exploration has been cast as a special case of convex RL (Hazan et al., 2019; Zhang et al., 2020; Zahavy et al., 2023). In such framework, the general task is defined via an F-bounded concave [^2] In practice, the function can be either convex, concave, or even non-convex and the term is used to distinguish the objective from the standard (linear) RL objective. In the following, we will assume F is concave if not mentioned otherwise. utility function $\mathcal{F}:\Delta_{\mathcal{S}}\rightarrow(-\infty,F]$, with $F<\infty$, that is a function of the state distribution $d^{\pi}$. This allows for a generalization of the standard RL learning objective, which is linear with respect to the state distribution. Usually, some regularity assumptions are enforced on the function $\mathcal{F}$, the most common being:
Assumption 2.1 (Lipschitz).
A function $\mathcal{F}:\mathcal{A}\rightarrow\mathbb{R}$ is Lipschitz-continuous for some constant $L<\infty$, or L-Lipschitz for short, if it holds
$|\mathcal{F}(x)-\mathcal{F}(y)|\leq L\|x-y\|_{1},\;\forall(x,y)\in\mathcal{A}^{2}$
More recently, Mutti et al. (2023b) noticed that in many practical scenarios only a finite number of $K\in\mathbb{N}^{+}$ episodes/trials can be drawn while interacting with the environment, and in such cases one should focus on $d_{K}$ rather than $d^{\pi}$. As a result, they contrast the infinite-trials objective defined as $\zeta_{\infty}(\pi):=\mathcal{F}(d^{\pi})$ with a finite-trials one, namely $\zeta_{K}(\pi):=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi}_{K}}\mathcal{F}(d_{k})$, noticing that Convex MDPs are characterized by the fact that $\zeta_{K}(\pi)\leq\zeta_{\infty}(\pi)$, differently from standard (linear) MDPs for which equality holds. In general, task-agnostic exploration can be easily defined as solving a Convex MDP equipped with an entropy functional (Hazan et al., 2019), namely $\mathcal{F}(d^{\pi}):=H(d^{\pi})$.
@@ 3 Problem Formulation
(i) How can task-agnostic exploration be defined in MARL?
In fact, when a reward function is not available, the core of the problem resides in finding a well-behaved problem formulation coherent with the task. We start by introducing a general framework that is a convex generalization of MGs, namely a tuple $\mathcal{M}_{\mathcal{F}}:=(\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{F},\mu,T)$, that consists in a MG equipped with a (non-linear) function $\mathcal{F}(\cdot)$. We refer to these objects as a Convex Markov Games (CMGs). How much should the agents coordinate? How much information should they have access to? Different answers depict different objectives.
Joint Objectives.  The first and most straightforward way to formulate the problem is to define it as in the MDP setting, with the joint state distribution simply taking the place of the single-agent state distribution. In this case, we define a Joint objective, consisting of
$\vspace{-5pt}\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\zeta_{\infty}(\pi)$ $:=\mathcal{F}(d^{\pi})\Big{\}}$ (1)
$\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\zeta_{K}(\pi)$ $:=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi}_{K}}\mathcal{F}(d_{K})\Big{\}}$ (2)
In task-agnostic exploration tasks, i.e. by setting $\mathcal{F}(\cdot):=H(\cdot)$, an optimal (joint) policy will try to cover the joint state space as uniformly as possible, either in expectation or over a finite number of trials respectively. In this, the joint formulation is rather intuitive as it describes the most general case of multi-agent exploration. Moreover, as each agent sees a difference in performance explicitly linked to others, this objective should be able to foster coordinated exploration. As we will see, this comes at a price.
Disjoint Objectives.  One might look for formulations more coherent with a multi-agent setting. The most trivial option is to design a disjoint counterpart of the objectives, that means to define a set of functions supported on per-agent state distributions rather than joint distributions. This intuition leads to Disjoint objectives:
$\Big{\{}\max_{\pi^{i}\in\Pi^{i}}\zeta^{i}_{\infty}(\pi^{i},\pi^{-i})$ $:=\mathcal{F}(d_{i}^{\pi^{i},\pi^{-i}})\Big{\}}_{i\in[\mathcal{N}]}$ (3)
$\Big{\{}\max_{\pi^{i}\in\Pi^{i}}\zeta^{i}_{K}(\pi^{i},\pi^{-i})$ $:=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi^{i},\pi^{-i}}_{K}}\mathcal{F}(d_{K,i})\Big{\}}_{i\in[\mathcal{N}]}$ (4)
@@
Mixture Objectives.
At last, we introduce a problem formulation that will be later prove capable of uniquely taking advantage of the structure of the problem. In order to do so, we first introduce the following:
@@
The agents have the same state spaces, namely $\mathcal{S}_{i}=\mathcal{S}_{j}=\tilde{\mathcal{S}},\;\forall(i,j)\in\mathcal{N}\times\mathcal{N}$.  [^3] One should notice that even in CMGs where this is not (even partially) the case, the assumption can be enforced by padding together the per-agent states.
Under this assumption, now on we will drop the agent subscript when referring to the per-agent states, and use $\tilde{\mathcal{S}}$ instead. Interestingly, this assumptions allows us to define a particular distribution, namely:
$\tilde{d}^{\pi}(\tilde{s}):=\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}d^{\pi}_{i}(\tilde{s})\in\Delta_{\tilde{\mathcal{S}}}.$
We refer to this distribution as mixture distribution, given that it is defined as a uniform mixture of the per-agent marginal distributions. Intuitively, it describes the average probability over all the agents to be in a common state $\tilde{s}\in\tilde{\mathcal{S}}$, in contrast with the joint distribution that describes the probability for them to be in a joint state $s$, or the marginals that describes the probability of each one of them separately. In Figure 1 we provide a visual representation of these concepts.
Figure 1: The interaction on the left induces different (empirical) distributions: marginal distributions for agent 1 and agent 2 over their respective states; a joint distribution over the product space; a mixture distribution over a common space, defined as the average. The mixture distribution is usually less sparse.
@@
$\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\tilde{\zeta}_{\infty}(\pi)$ $:=\mathcal{F}(\tilde{d}^{\pi})\Big{\}}$ (5)
$\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\tilde{\zeta}_{K}(\pi)$ $:=\operatorname*{\mathbb{E}}_{\tilde{d}_{K}\sim p^{\pi}_{K}}\mathcal{F}(\tilde{d}_{K})\Big{\}}$ (6)
By employing this kind of objectives in the task-agnostic exploration, i.e. by setting $\mathcal{F}(\cdot):=H(\cdot)$, two are the possible scenarios of optimal behaviors. In the first scenario, each agent tries to cover her state space as uniformly as possible, without taking into account the presence of others. In this sense, the mixture objectives enforce a behavior similar to the disjoint ones. The second scenario is more interesting and it has been referred to in Kolchinsky & Tracey (2017) for general distributions as the clustered scenario: agents will form a grouping such that marginal distributions in the same group are approximately the same, while distributions assigned to different groups will be very different from one another, potentially with a disjoint support. In other words, agents will try to cover different sub-portions of the state space in groups, so that, on average, all the state space will be covered uniformly. This second scenario is of particular interest for task-agnostic exploration, as it is the only one among the one presented that explicitly enforces policies inducing diverse state distributions among the agents.
Remarks.
Mixture objectives in task-agnostic exploration require estimating the entropy of mixture state-distributions, which might remain challenging in high-dimensional scenarios. Fortunately, the problem of efficiently estimating the entropy for general mixture distributions has been previously investigated in Kolchinsky & Tracey (2017) and extended to RL with mixture policies by Baram et al. (2021). The same ideas might be applied to the case of interest.

### v3

# Towards Unsupervised Multi-Agent Reinforcement Learning via Task-Agnostic Exploration

@@ Abstract
In reinforcement learning, we typically refer to unsupervised pre-training when we aim to pre-train a policy without a priori access to the task specification, i.e., rewards, to be later employed for efficient learning of downstream tasks. In single-agent settings, the problem has been extensively studied and mostly understood. A popular approach, called task-agnostic exploration, casts the unsupervised objective as maximizing the entropy of the state distribution induced by the agent’s policy, from which principles and methods follow. In contrast, little is known about it in multi-agent settings, which are ubiquitous in the real world. What are the pros and cons of alternative problem formulations in this setting? How hard is the problem in theory, how can we solve it in practice? In this paper, we address these questions by first characterizing those alternative formulations and highlighting how the problem, even when tractable in theory, is non-trivial in practice. Then, we present a scalable, decentralized, trust-region policy search algorithm to address the problem in practical settings. Finally, we provide numerical validations to both corroborate the theoretical findings and pave the way for unsupervised multi-agent reinforcement learning via task-agnostic exploration in challenging domains, showing that optimizing for a specific objective, namely mixture entropy, provides an excellent trade-off between tractability and performances.
@@ 1 Introduction
Multi-Agent Reinforcement Learning [MARL, 2] recently showed promising results in learning complex behaviors, such as coordination and teamwork [35], strategic planning in the presence of imperfect knowledge [33], and trading [11]. Just like in single-agent RL, however, most of the efforts are focused on tabula rasa learning, that is, without exploiting any prior knowledge gathered from offline data and/or policy pre-training. Despite its generality, learning tabula rasa hinders MARL from addressing real-world situations, where training from scratch is slow, expensive, and arguably unnecessary [1]. In this regard, some progress has been made on techniques specific to the multi-agent setting, ranging from ad hoc teamwork [23] to zero-shot coordination [9], but our understanding of what can be done instead of learning tabula rasa is still limited.
@@
In single-agent RL, unsupervised pre-training frameworks [13] have emerged as a viable solution: a policy is pre-trained without a priori access to the task specification, i.e., rewards, to be later employed for efficient learning of downstream tasks. Among others, state-entropy maximization [8, 14] was shown to be a useful tool for policy pre-training [8, 25] and data collection for offline learning [44]. In this setting, the unsupervised objective is casted as maximizing the entropy of the state distribution induced by the agent’s policy. Recently, the potential of entropy objectives in MARL was empirically corroborated by a plethora of works [19, 50, 42, 41] investigating entropic reward-shaping techniques to boost exploration in downstream tasks. Yet, to the best of our knowledge, we are still lacking a principled understanding of how task-agnostic exploration works in multi-agent settings, and how it can be used as an unsupervised pre-training fashion. Let us think of an illustrative example that highlights the central question of this work: multiple autonomous robots deployed in a factory for a production task. The robots’ main goal is to perform many tasks over a large set of products, with objectives such as optimizing for costs or energy or throughput might be changing over time depending on the market’s condition. Arguably, trying to learn each possible task from scratch is inefficient and unnecessary. On the other hand, one could think to first learn to cover the possible states of the system, to learn its dynamics, and then fine-tune over a specific task. Yet, if everyone is focused on their own exploration, any incentive to collaborate with each other may disappear, especially when coordinating comes at a cost. Similarly, covering the entire space might be unreasonable in most real-world cases. Clearly, a third option is needed.
Research Questions and Paper Structure.  Throughout this paper, we will address the following questions: (Q1) Can we formulate in a principled way unsupervised pre-training via task-agnostic exploration in MARL as well? (Section 3);(Q2) How are different formulations related? Do crucial theoretical differences emerge? (Section 4); (Q3) Can we explicitly pre-train a policy for task-agnostic exploration in practical multi-agent scenarios? (Section 5); (Q4) Do crucial differences emerge in practice? Does this have an impact on downstream tasks learning? (Section 6).
@@ 2 Preliminaries
Notation.  We denote $[N]:=\{1,2,\ldots,N\}$ for a constant $N<\infty$. We denote a set with a calligraphic letter $\mathcal{A}$ and its size as $|\mathcal{A}|$. For a (finite) set $\mathcal{A}=\{1,2,\dots,i,\dots\}$, we denote $-i=\mathcal{A}/\{i\}$ the set of all its elements but $i$. $\mathcal{A}^{T}:=\times_{t=1}^{T}\mathcal{A}$ is the $T$-fold Cartesian product of $\mathcal{A}$. The simplex on $\mathcal{A}$ is $\Delta_{\mathcal{A}}:=\{p\in[0,1]^{|\mathcal{A}|}|\sum_{a\in\mathcal{A}}p(a)=1\}$ and $\Delta_{\mathcal{A}}^{\mathcal{B}}$ denotes the set of conditional distributions $p:\mathcal{A}\to\Delta_{\mathcal{B}}$. Let $X,X^{\prime}$ random variables on the set of outcomes $\mathcal{X}$ and corresponding probability measures $p_{X},p_{X^{\prime}}$, we denote the Shannon entropy of $X$ as $H(X)=-\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x))$ and the Kullback-Leibler (KL) divergence as $D_{\mathrm{KL}}(p_{X}\|p_{X^{\prime}})=\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x)/p_{X^{\prime}}(x))$. We denote $\mathbf{x}=(X_{1},\ldots,X_{T})$ a random vector of size $T$ and $\mathbf{x}[t]$ its entry at position $t\in[T]$.
Interaction Protocol.  As a model for interaction, we consider finite-horizon Markov Games [MGs, 16] without rewards. A MG $\mathcal{M}:=(\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},\mu,T)$ is composed of a set of agents $\mathcal{N}$, a set $\mathcal{S}=\times_{i\in[\mathcal{N}]}\mathcal{S}_{i}$ of states, and a set of (joint) actions $\mathcal{A}=\times_{i\in[\mathcal{N}]}\mathcal{A}_{i}$, which we let discrete and finite with size $|\mathcal{S}|,|\mathcal{A}|$ respectively. At the start of an episode, the initial state $s_{1}$ of $\mathcal{M}$ is drawn from an initial state distribution $\mu\in\Delta_{\mathcal{S}}$. Upon observing $s_{1}$, each agent takes action $a_{1}^{i}\in\mathcal{A}_{i}$, the system transitions to $s_{2}\sim\mathbb{P}(\cdot|s_{1},a_{1})$ according to the transition model $\mathbb{P}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$. The process is repeated until $s_{T}$ is reached and $s_{T}$ is generated, being $T<\infty$ the horizon of an episode. Each agent acts according to a policy, that can be either Markovian, i.e. $\pi^{i}\in\Delta_{\mathcal{S}}^{\mathcal{A}^{i}}$, or Non-Markovian over, i.e. $\pi^{i}\in\Delta_{\mathcal{S}^{t}\times\mathcal{A}^{t}}^{\mathcal{A}^{i}}$. [^1] In general, we will denote the set of valid per-agent policies with $\Pi^{i}$ and the set of joint policies with $\Pi$. Also, we will denote as decentralized-information policies the ones conditioned on either $\mathcal{S}_{i}$ or $\mathcal{S}^{t}_{i}\times\mathcal{A}^{t}_{i}$ for agent $i$, and centralized-information ones the ones conditioned over the full state or state-actions sequences. It follows that the joint action is taken according to the joint policy $\Delta_{\mathcal{S}}^{\mathcal{A}}\ni\pi=(\pi^{i})_{i\in[\mathcal{N}]}$.
@@
Convex MDPs and Task-Agnostic Exploration.  In the MDP setting ($|\mathcal{N}|=1$), the problem of task-agnostic exploration, together with others as imitation-learning, diverse skill discovery and inverse RL, is a special case of convex RL [8, 48, 45]. In such framework, the general task is defined via an F-bounded concave [^2] In practice, the function can be either convex, concave, or even non-convex. The term is used to distinguish the objective from the standard (linear) RL objective. We will assume $\mathcal{F}$ is concave if not mentioned otherwise. utility function $\mathcal{F}:\Delta_{\mathcal{S}}\rightarrow(-\infty,F]$, with $F<\infty$, that is a function of the state distribution $d^{\pi}$. This allows for a generalization of the standard RL objective, which is a linear product between a reward vector and the state(-action) distribution [34]. Usually, some regularity assumptions are enforced on the function $\mathcal{F}$. In the following, we align with the most common Lipschitz assumption, explicitly described in Appendix B. More recently, [26] noticed that in many practical scenarios only a finite number of $K\in\mathbb{N}^{+}$ episodes/trials can be drawn while interacting with the environment, and in such cases one should focus on $d_{K}$ rather than $d^{\pi}$. As a result, they contrast the infinite-trials objective defined as $\zeta_{\infty}(\pi):=\mathcal{F}(d^{\pi})$ with a finite-trials one, namely $\zeta_{K}(\pi):=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi}_{K}}\mathcal{F}(d_{k})$, noticing that convex MDPs (cMDPs) are characterized by the fact that $\zeta_{K}(\pi)\leq\zeta_{\infty}(\pi)$, differently from standard (linear) MDPs for which equality holds. In single-agent convex RL, task-agnostic exploration is defined as solving a cMDP equipped with an entropy functional [8], namely $\mathcal{F}(d^{\pi}):=H(d^{\pi})$.
@@ 3 Problem Formulation
(Q1) Can we formulate in a principled way unsupervised pre-training
via task-agnostic exploration in MARL as well?
In fact, when a reward function is not available, the core of the problem resides in finding a well-behaved problem formulation coherent with the task. [7] recently introduced a convex generalization of MGs called convex Markov Games (cMGs), namely a tuple $\mathcal{M}_{\mathcal{F}}:=(\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{F},\mu,T)$, that consists in a MG equipped with (non-linear) functions of the stationary joint state distribution $\mathcal{F}(d^{\pi})$. We expand over this definition, by noticing that task-agnostic exploration can be casted as solving a cMG equipped with an entropy functional, namely $\mathcal{F}(\cdot):=H(\cdot)$. Yet, important new questions arise: Over which distributions should agents compute the entropy? How much information should they have access to? Can we define objectives accounting for a finite number of trials? Different answers depict different objectives.
Joint Objectives.  The first and most straightforward way to formulate the problem is to define it as in the MDP setting, with the joint state distribution simply taking the place of the single-agent state distribution. In this case, we define infinite-trials and finite-trials Joint objectives, respectively
$\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\zeta_{\infty}(\pi):=\mathcal{F}(d^{\pi})\Big{\}}$ $\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\zeta_{K}(\pi):=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi}_{K}}\mathcal{F}(d_{K})\Big{\}}$ (1)
In task-agnostic exploration tasks, an optimal (joint) policy will try to cover the joint state space uniformly, either in expectation or over a finite number of trials respectively. In this, the joint formulation is rather intuitive as it describes the most general case of multi-agent exploration. Moreover, as each agent sees a difference in performance explicitly linked to others, this objective should be able to foster coordinated exploration. As we shall see, this comes at a price.
Disjoint Objectives.  One might look for formulations that fully embrace the multi-agent setting, such as defining a set of functions supported on per-agent state distributions rather than joint distributions. This intuition leads to infinite-trials and finite-trials Disjoint objectives:
$\Big{\{}\max_{\pi^{i}\in\Pi^{i}}\zeta^{i}_{\infty}(\pi^{i},\cdot):=\mathcal{F}(d_{i}^{\pi^{i},\cdot})\Big{\}}_{i\in[\mathcal{N}]}$$\Big{\{}\max_{\pi^{i}\in\Pi^{i}}\zeta^{i}_{K}(\pi^{i},\cdot):=\operatorname*{\mathbb{E}}_{d_{K}\sim p^{\pi^{i},\cdot}_{K}}\mathcal{F}(d_{K,i})\Big{\}}_{i\in[\mathcal{N}]}$ (2)
@@
Mixture Objectives.  At last, we introduce a problem formulation that will later prove capable of uniquely taking advantage of the structure of the problem. First, we introduce the following:
@@
The agents have the same state space $\mathcal{S}_{i}=\mathcal{S}_{j}=\tilde{\mathcal{S}},\forall(i,j)\in\mathcal{N}\times\mathcal{N}$ [^3] One should notice that even in cMGs where this is not (even partially) the case, the assumption can be enforced by padding together the per-agent states.
Under this assumption, we will drop the agent subscript when referring to the per-agent states and use $\tilde{\mathcal{S}}$ instead. Interestingly, this assumption allows us to define a particular distribution: $\tilde{d}^{\pi}(\tilde{s}):=\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}d^{\pi}_{i}(\tilde{s})\in\Delta_{\tilde{\mathcal{S}}}.$ We refer to this distribution as mixture distribution, given that it is defined as a uniform mixture of the per-agent marginal distributions. Intuitively, it describes the average probability over all the agents to be in a common state $\tilde{s}\in\tilde{\mathcal{S}}$, in contrast with the joint distribution that describes the probability for them to be in a joint state $s$, or the marginals that describes the probability of each one of them separately.
@@
$\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\tilde{\zeta}_{\infty}(\pi):=\mathcal{F}(\tilde{d}^{\pi})\Big{\}}$ $\max_{\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}}\ \Big{\{}\tilde{\zeta}_{K}(\pi):=\operatorname*{\mathbb{E}}_{\tilde{d}_{K}\sim p^{\pi}_{K}}\mathcal{F}(\tilde{d}_{K})\Big{\}}$ (3)
When this kind of objectives is employed in task-agnostic exploration, the entropy of the mixture distribution decomposes as $H(\tilde{d}^{\pi})=\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}H(d^{\pi}_{i})+\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}D_{\mathrm{KL}}(d^{\pi}_{i}||\tilde{d}^{\pi})$ and one remarkable scenario arises: Agents follow policies possibly inducing lower disjoint entropies, but their induced marginal distributions are maximally different. Thus, the average entropy remains low, but the overall mixture entropy is high due to diversity (i.e., high values of the KL divergences). This scenario has been referred to in [12] as the clustering scenario and, in the following, we will provide additional evidences why this scenario is particularly relevant.